OpenStudy (anonymous):

Calculate the distance between the point (1,8) and the line l: -3x + y = -5.

OpenStudy (anonymous):

u first

OpenStudy (anonymous):

There could many distance .. do you want the shortest ?

OpenStudy (anonymous):

Yes please!!

OpenStudy (anonymous):

It's simple then, the distance is given by $$\huge|\small\frac{(-3)(1)+8+5}{\sqrt{(-3)^2+(1)^2}}\huge|\small =\sqrt{10}$$

OpenStudy (anonymous):

If you want to understand why this holds : http://demonstrations.wolfram.com/ShortestDistanceBetweenAPointAndALineIn2D/